# On semidefinite descriptions for convex hulls of quadratic programs

### 07 Mar 2024 v1

@@ 3 Convex hull exactness
This section presents necessary and sufficient conditions for convex hull exactness, i.e., the property that $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$. The following result rephrases convex hull exactness as a question regarding the existence of nontrivial “rounding directions.” This result follows from basic convex analysis and is proved in Appendix A.
@@
${\cal R}(x,t)\coloneqq\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\exists\epsilon>0\text{ s.t. }[(x,t)\pm\epsilon(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}\right\}.$
We say that this set is nontrivial if it contains a nonzero element.
Note that ${\cal R}(x,t)$ is a convex cone satisfying $-{\cal R}(x,t)={\cal R}(x,t)$ so that ${\cal R}(x,t)$ is indeed a subspace.
Remark 3.
One may compare our rounding directions to other similar definitions from elementary convex analysis [11, Section 5.1]. Fix a point $(x,t)\in{\cal S}_{\textup{SDP}}$ and $(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}$. Recall that $(x^{\prime},t^{\prime})$ is a feasible direction if there exists $\alpha>0$ such that $[(x,t),(x,t)+\alpha(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}$. In particular, feasible directions are a unidirectional notion, whereas rounding directions are bidirectional. Next, recall that $(x^{\prime},t^{\prime})$ is a tangent direction if it is a limit of feasible directions. Again, tangent directions are unidirectional.
@@ 3.1 Sufficient conditions for convex hull exactness
In general, the subspace of rounding directions ${\cal R}(x,t)$ is a difficult object to describe explicitly. Below, we identify a more manageable subset of ${\cal R}(x,t)$ that we will refer to as ${\cal R}^{\prime}(x,t)$. This will lead to a sufficient condition for convex hull exactness upon invoking Lemma 1.
@@
If we assume (incorrectly) that $q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}$ is linear in $\alpha$, then as ${\cal G}(x,t)$ is the minimal face of $\Gamma^{\circ}$ containing $q(x)-2te_{\textup{obj}}$, we have $q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\in\operatorname{span}({\cal G}(x,t))$ for all $\alpha\in[-\epsilon,\epsilon]$ (in fact, also for all $\alpha\in{\mathbb{R}}$). Thus, $(x^{\prime},t^{\prime})\in{\cal R}^{\prime}(x,t)$. Unfortunately, $q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}$ is nonlinear in $\alpha$ so that not all rounding directions from ${\cal R}(x,t)$ can be captured in this form. Below, we show that ${\cal R}^{\prime}(x,t)\subseteq{\cal R}(x,t)$. We will later give additional intuition and motivation for ${\cal R}^{\prime}(x,t)$ when $\Gamma^{\circ}$ is facially exposed.
@@
The condition of Theorem 1 on the surface requires checking infinitely many systems of infinitely many constraints. Nonetheless, it is possible to check these conditions in some situations where $\Gamma$ (and related sets) possess simplifying symmetries. See Appendix B for some worked examples.
@@ 3.2 Necessary conditions for convex hull exactness
We now show that under a technical assumption (Assumption 2), we have ${\cal R}^{\prime}(x,t)={\cal R}(x,t)$. Thus, the sufficient condition of Theorem 1 is also necessary under the same assumption.
@@
This assumption holds for any cone isomorphic to a slice of the nonnegative orthant, the second-order cone, or the positive semidefinite cone. See [19] for a longer discussion of this assumption and its connections to the nice cones. In general, all nice cones are facially exposed.
We will use the following property of exposed faces ${\cal G}\trianglelefteq\Gamma^{\circ}$.
@@
As each coordinate of this vector is quadratic in $\alpha$, it suffices to show instead that the above inclusion holds for all $\alpha\in[-1,1]$. Indeed, suppose this inclusion has been shown for all $\alpha\in[-1,1]$ and let $\ell\in\operatorname{span}({\cal G}(x,t))^{\perp}$. Then, $\left\langle\ell,q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\right\rangle$ is a quadratic function in $\alpha$ mapping the interval $[-1,1]$ to 00. In particular, it must map all of ${\mathbb{R}}$ to zero. As $\ell\in\operatorname{span}({\cal G}(x,t))^{\perp}$ was arbitrary, we deduce that $q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\in\operatorname{span}({\cal G}(x,t))$ for all $\alpha\in{\mathbb{R}}$.
@@ 3.3 Explicit descriptions of ${\cal R}^{\prime}(x,t)$
Below, we give slightly simplified descriptions of ${\cal R}^{\prime}(x,t)$ (see Propositions 1 and 2) that may be useful in verifying the conditions of Theorems 1 and 2. See, for example, their use in Appendix B. The derivations of these descriptions require the following observation.
@@
The proof of 1 and of Propositions 1 and 2 below are given in Appendix A.
@@
Suppose Assumption 1 holds and let $(x,t)\in{\cal S}_{\textup{SDP}}$ where $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$. Then,
${\cal R}^{\prime}(x,t)=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}(x^{\prime})^{\top}A[\gamma]x^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\\ \left\langle A[\gamma]x+b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\end{array}\right\}.$
Proposition 2.
Suppose Assumptions 1 and 2 hold. Let $(x,t)\in{\cal S}_{\textup{SDP}}$ where $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$ and let $(1,f)\in\operatorname{rint}({\cal F}(x,t))$. Then,
${\cal R}^{\prime}(x,t)=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime}\in\ker(A[f]),\\ \left\langle A[\eta]x+b[\eta],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\eta)\in{\cal G}(x,t)^{\perp}\end{array}\right\}.$
3.4 Revisiting the setting of polyhedral $\Gamma$
Wang and Kılınç-Karzan [25] give sufficient conditions for convex hull exactness under the assumption that $\Gamma_{1}$ is polyhedral (equivalently, by Remark 1, whenever $\Gamma$ is polyhedral). This assumption holds, for example, when the set of quadratic forms $\left\{A_{\textup{obj}},A_{1},\dots,A_{m}\right\}$ is simultaneously diagonalizable. Specializing Theorem 2 to this setting, we prove the following necessary and sufficient counterpart to [25, Theorem 1].
{restatable}
${\cal R}^{\prime}(x,t)=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime}\in\ker(A[f]),\\ \left\langle b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal F}(x,t)\end{array}\right\}.$
Note that the only dependence on $(x,t)$ in the above description of ${\cal R}^{\prime}$ is in the definition of ${\cal F}={\cal F}(x,t)$. Hence, $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$ if and only if the above set is nontrivial for every ${\cal F}\trianglelefteq\Gamma$ which is exposed by some vector $q(x)-2te_{\textup{obj}}$ for $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$.
@@ 3.4 Revisiting the setting of polyhedral $\Gamma$
We begin by noting that when $\Gamma$ is polyhedral, so too is $\Gamma^{\circ}$ so that Assumption 2 holds. Next, we claim that for every face ${\cal G}\trianglelefteq\Gamma^{\circ}$ we have ${\cal G}^{\perp}=\operatorname{span}({\cal G}^{\triangle})$. By definition, $\operatorname{span}({\cal G}^{\triangle})=\operatorname{span}(\Gamma\cap{\cal G}^{\perp})\subseteq{\cal G}^{\perp}$. On the other hand, as $\Gamma$ and $\Gamma^{\circ}$ are polyhedral, we have that $\dim({\cal G})+\dim({\cal G}^{\triangle})=m$ [22, Theorem 3]. Rearranging this equation, we have $\dim({\cal G}^{\triangle})=m-\dim({\cal G})=\dim({\cal G}^{\perp})$. We conclude that ${\cal G}^{\perp}=\operatorname{span}({\cal G}^{\triangle})$.
Here, the second line follows because we have shown ${\cal G}^{\perp}=\operatorname{span}({\cal G}^{\triangle})$ holds for every face ${\cal G}\trianglelefteq\Gamma^{\circ}$ and by definition ${\cal F}(x,t)={\cal G}(x,t)^{\triangle}$. The third line follows from the fact that $(1,f)\in\operatorname{rint}({\cal F}(x,t))$ implies $\ker(A[f])\subseteq\ker(A[\gamma])$ for every $(1,\gamma)\in{\cal F}(x,t)$. The result then follows from Theorem 2. ∎
Remark 5.
The main difference between Section 3.4 and [25, Theorem 1] is that Section 3.4 only considers certain (a fortiori semidefinite) faces of $\Gamma$ whereas [25, Theorem 1] imposes a constraint on every semidefinite face of $\Gamma$. This idea of restricting the analysis to certain faces of $\Gamma$ was previously investigated by [6, 17] who used it to provide sufficient conditions for objective value exactness.
Section 3.4 can be used to characterize when the SDP relaxation of a set defined by two quadratic constraints is exact; see Appendix B. This complements work of [26] that describes the convex hull of a set defined by two quadratic constraints in all settings but requires a strengthened SDP relaxation.
Acknowledgments
This work was supported by the National Science Foundation [grant number CMMI 1454548], the Office of Naval Research [grant number N00014-19-1-2321], and the Air Force Office of Scientific Research [grant number FA95502210365].
Appendix A Deferred proofs
Proof of Remark 1.
Fix $(x,t)\in{\mathbb{R}}^{n+1}$. Note that
$\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$ $=\sup_{\gamma\in{\mathbb{R}}_{+}^{m}}\left\{[\gamma,q(x)]:\,A[\gamma]\succeq 0\right\}$
$=\inf_{\xi\in{\mathbb{S}}^{n}_{+}}\left\{q_{\textup{obj}}(x)+\left\langle A_{\textup{obj}},\xi\right\rangle:\,q_{i}(x)+\left\langle A_{i},\xi\right\rangle\leq 0,\forall i\in[m]\right\}.$
Proof of Lemma 1.
Note that ${\cal S}_{\textup{SDP}}$ is a closed convex set containing no lines. It is clear that $(0_{n},1)$ is a recessive direction of ${\cal S}_{\textup{SDP}}$. We claim that $(0_{n},1)$ is the only recessive direction: Let $\gamma^{*}$ such that $A[\gamma^{*}]\succ 0$ (this is possible by Assumption 1) and consider any $(x^{\prime},t^{\prime})$ where $x^{\prime}$ is nonzero. Then, for any $(\tilde{x},\tilde{t})\in{\cal S}_{\textup{SDP}}$, we have $2(\tilde{t}+\alpha t^{\prime})<[\gamma^{*},q(\tilde{x}+\alpha x^{\prime})]$ for all $\alpha>0$ large enough. We deduce by [20, Theorem 18.5], that ${\cal S}_{\textup{SDP}}$ is the sum of the convex hull of its extreme points and the direction $(0_{n},1)$. In particular $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$ if and only if $(x,t)$ is not extreme for every $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$. By definition, $(x,t)$ is not extreme if and only if there exists a nonzero $(x^{\prime},t^{\prime})$ and $\alpha>0$ such that $[(x,t)\pm\alpha(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}$. ∎
Proof of 1.
Suppose $\operatorname{span}({\cal G}(x,t))\supseteq{\mathbb{R}}\times 0_{m}$ so that $(0_{n},1)\in{\cal R}^{\prime}(x,t)$. By Lemma 2, there exists $\alpha>0$ such that $(x,t-\alpha)\in{\cal S}_{\textup{SDP}}$. This contradicts $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$. Thus, we deduce that $\operatorname{span}({\cal G}(x,t))\not\supseteq{\mathbb{R}}\times 0_{m}$. Equivalently, ${\cal G}(x,t)^{\perp}\not\subseteq 0\times{\mathbb{R}}^{m}$ and there exists $(1,\bar{\gamma})\in{\cal G}(x,t)^{\perp}$. Then, for any $(\gamma_{\textup{obj}},\gamma)\in{\cal G}(x,t)^{\perp}$, we can write $(\gamma_{\textup{obj}},\gamma)$ as a linear combination of $(\gamma_{\textup{obj}},\gamma)+(1-\gamma_{\textup{obj}})(1,\bar{\gamma})$ and $(1,\bar{\gamma})$. ∎
Proof of Proposition 1.
@@ Appendix A Deferred proofs
Proof of Proposition 2.
Let $(1,f)\in\operatorname{rint}({\cal F}(x,t))$. By Proposition 1, it suffices to show that $x^{\prime}\in\ker(A[f])$ if and only if
@@
Appendix B Example applications of Theorems 1, 2 and 3.4
In this section, we apply the results of Section 3 to three problems. These examples show that the conditions of Theorems 1, 2 and 3.4 can be checked in situations where $\Gamma$ is sufficiently simple.
B.1 Mixed binary programming
To begin, we apply our results to a well-studied prototypical set involving a convex quadratic function, a binary variable and a big-M relation. This set occurs as a substructure commonly studied in sparse regression applications. The convex hull characterization of this set is well-known in the literature and is often shown as a consequence of the perspective formulation trick due to Ceria and Soares [7]. The example in this subsection highlights the systematic nature of our approach.
Consider the epigraph set
${\cal S}=\left\{(x,t)\in{\mathbb{R}}^{2}\times{\mathbb{R}}:\,\begin{array}[]{l}q_{\textup{obj}}(x)\coloneqq x_{2}^{2}\leq 2t\\ q_{1}(x)\coloneqq x_{1}(x_{1}-1)=0\\ q_{2}(x)\coloneqq\sqrt{2}x_{2}(x_{1}-1)=0\end{array}\right\}.$
In words, $x_{1}$ is a binary on-off variable, $x_{2}$ is a continuous variable which is constrained to be off whenever $x_{1}$ is off, and $t$ is the epigraph variable corresponding to $x_{2}^{2}$. The normalization of $q_{2}(x)$ is not important here and is made only for notational convenience in the calculations. Additionally, the presence of equality constraints does not markedly affect our results; the only change in this case is that the corresponding $\gamma$ multiplier is no longer constrained to be nonnegative.
It is well-known that $\operatorname{conv}({\cal S})$ is given by the perspective reformulation of ${\cal S}$ (see e.g., [9]), i.e.,
[…]
We give an alternative proof of (6). We will show that $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$, the projected SDP relaxation, using Theorem 2. Then, using an explicit description of $\Gamma^{\circ}$, we will give a description of $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$ in the original space.
A simple computation shows that in this setting, we have
$\Gamma=\left\{(\gamma_{\textup{obj}},\gamma)\in{\mathbb{R}}^{3}:\,\gamma_{\textup{obj}}+\gamma_{1}\geq\sqrt{(\gamma_{\textup{obj}}-\gamma_{1})^{2}+\left(\sqrt{2}\gamma_{2}\right)^{2}}\right\}\text{ and}$
$\Gamma^{\circ}=\left\{(\ell_{\textup{obj}},\ell)\in{\mathbb{R}}^{3}:\,-\ell_{\textup{obj}}-\ell_{1}\geq\sqrt{(\ell_{\textup{obj}}-\ell_{1})^{2}+\left(\sqrt{2}\ell_{2}\right)^{2}}\right\}.$
In words, $\Gamma$ and $\Gamma^{\circ}$ are both (rotated) second-order cones and Assumptions 1 and 2 hold.
It remains to show that for all $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$, the set ${\cal R}^{\prime}(x,t)$ is nontrivial. To this end, let $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$. Note that as the point $(0,0)\in{\cal S}$, we deduce $(x,t)\neq(0,0)$. Recall that $\Gamma^{\circ}$ has three types of faces: the two trivial faces (the apex and the cone itself) and the one-dimensional proper faces. Thus, there are three cases to consider: (i) ${\cal G}(x,t)=\left\{0\right\}$, (ii) ${\cal G}(x,t)=\Gamma^{\circ}$, and (iii) ${\cal G}(x,t)$ is a one-dimensional face of $\Gamma^{\circ}$.
In case (i), $q(x)-2te_{\textup{obj}}=0$ implying that $(x,t)\in{\cal S}$, a contradiction. In case (ii), $\operatorname{span}({\cal G}(x,t))={\mathbb{R}}^{3}$ so that ${\cal R}^{\prime}(x,t)={\mathbb{R}}^{3}$ and is nontrivial. In the final case, a mechanical application of Proposition 2 gives

### 20 Mar 2024 v2

@@ 3 Convex hull exactness
This section presents necessary and sufficient conditions for convex hull exactness, i.e., the property that $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$. Below, we rephrase convex hull exactness as a question regarding the existence of nontrivial “rounding directions.”
@@
Proof.
Note that ${\cal S}_{\textup{SDP}}$ is a closed convex set containing no lines. It is clear that $(0_{n},1)$ is a recessive direction of ${\cal S}_{\textup{SDP}}$. We claim that $(0_{n},1)$ is the only recessive direction: Let $\gamma^{*}$ such that $A[\gamma^{*}]\succ 0$ (this is possible by Assumption 1) and consider any $(x^{\prime},t^{\prime})$ where $x^{\prime}$ is nonzero. Then, for any $(\tilde{x},\tilde{t})\in{\cal S}_{\textup{SDP}}$, we have $2(\tilde{t}+\alpha t^{\prime})<[\gamma^{*},q(\tilde{x}+\alpha x^{\prime})]$ for all $\alpha>0$ large enough. We deduce by [15, Theorem 18.5], that ${\cal S}_{\textup{SDP}}$ is the sum of the convex hull of its extreme points and the direction $(0_{n},1)$. In particular $\operatorname{conv}({\cal S})={\cal S}_{\textup{SDP}}$ if and only if $(x,t)$ is not extreme for every $(x,t)\in{\cal S}_{\textup{SDP}}\setminus{\cal S}$. By definition, $(x,t)$ is not extreme if and only if there exists a nonzero $(x^{\prime},t^{\prime})$ and $\alpha>0$ such that $[(x,t)\pm\alpha(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}$. ∎
@@
${\cal R}(x,t)\coloneqq\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}\exists\epsilon>0:\\ {}[(x,t)\pm\epsilon(x^{\prime},t^{\prime})]\subseteq{\cal S}_{\textup{SDP}}\end{array}\right\}.$
This set is nontrivial if it is not $\left\{0\right\}$.
Note that ${\cal R}(x,t)$ is a convex cone and $-{\cal R}(x,t)={\cal R}(x,t)$ so ${\cal R}(x,t)$ is indeed a subspace.
@@ 3.1 Sufficient conditions for convex hull exactness
In general, ${\cal R}(x,t)$ is difficult to describe explicitly. Below, we identify a more manageable subset ${\cal R}^{\prime}(x,t)\subseteq{\cal R}(x,t)$. This will lead to a sufficient condition for convex hull exactness via lemma 1.
@@
If we assume (incorrectly) that $q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}$ is linear in $\alpha$, then this holds if and only if
$q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\in\operatorname{span}({\cal G}(x,t))$
for all $\alpha\in[-\epsilon,\epsilon]$ (in fact, also for all $\alpha\in{\mathbb{R}}$). This is precisely the definition of ${\cal R}^{\prime}(x,t)$. Thus, ${\cal R}^{\prime}(x,t)$ is a “first-order approximation” to ${\cal R}(x,t)$ in this sense. We will later give additional motivation for ${\cal R}^{\prime}(x,t)$ when $\Gamma^{\circ}$ is facially exposed.
Below, we show that ${\cal R}^{\prime}(x,t)\subseteq{\cal R}(x,t)$.
@@
The condition of theorem 1 on the surface requires checking infinitely many systems of infinitely many constraints. Nonetheless, it is possible to check these conditions in some situations where $\Gamma$ is sufficiently simple. See section 4.
@@ 3.2 Necessary conditions for convex hull exactness
We now show that, under a technical assumption (Assumption 2), ${\cal R}^{\prime}(x,t)={\cal R}(x,t)$ so that the sufficient condition of theorem 1 is also necessary.
@@
This assumption holds for any conic slice of the nonnegative orthant, the second-order cone, or the PSD cone. See [14] for a discussion of this assumption and its connections to the nice cones. In general, all nice cones are facially exposed.
We will use the following property of exposed faces.
@@
As each coordinate of this vector is quadratic in $\alpha$, it suffices to show instead that the inclusion holds for all $\alpha\in[-1,1]$. Indeed, suppose this inclusion has been shown for all $\alpha\in[-1,1]$ and let $\ell\in\operatorname{span}({\cal G}(x,t))^{\perp}$. Then, $\left\langle\ell,q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\right\rangle$ is a quadratic function in $\alpha$ mapping the interval $[-1,1]$ to 00. Thus, it must map all of ${\mathbb{R}}$ to zero. As $\ell\in\operatorname{span}({\cal G}(x,t))^{\perp}$ was arbitrary, we deduce that $q(x+\alpha x^{\prime})-2(t+\alpha t^{\prime})e_{\textup{obj}}\in\operatorname{span}({\cal G}(x,t))$ for all $\alpha\in{\mathbb{R}}$.
@@ 3.3 Explicit descriptions of ${\cal R}^{\prime}(x,t)$
propositions 1 and 2 below give alternate descriptions of ${\cal R}^{\prime}(x,t)$ (see propositions 1 and 2) that may be useful in verifying the conditions of theorems 1 and 2. See, for example, their use in section 4. We begin by making an observation and then proving propositions 1 and 2.
@@
Proof.
Suppose $\operatorname{span}({\cal G}(x,t))\supseteq{\mathbb{R}}\times 0_{m}$ so that $(0_{n},1)\in{\cal R}^{\prime}(x,t)$. By lemma 2, there exists $\alpha>0$ so that $(x,t-\alpha)\in{\cal S}_{\textup{SDP}}$. This contradicts $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$. Thus, $\operatorname{span}({\cal G}(x,t))\not\supseteq{\mathbb{R}}\times 0_{m}$. Equivalently, ${\cal G}(x,t)^{\perp}\not\subseteq 0\times{\mathbb{R}}^{m}$ and there exists $(1,\bar{\gamma})\in{\cal G}(x,t)^{\perp}$. Then, for any $(\gamma_{\textup{obj}},\gamma)\in{\cal G}(x,t)^{\perp}$, we can write $(\gamma_{\textup{obj}},\gamma)$ as a linear combination of $(\gamma_{\textup{obj}},\gamma)+(1-\gamma_{\textup{obj}})(1,\bar{\gamma})$ and $(1,\bar{\gamma})$. ∎
@@
Suppose Assumption 1 holds and $(x,t)\in{\cal S}_{\textup{SDP}}$ with $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$. Then,
${\cal R}^{\prime}(x,t)$ =$=\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}(x^{\prime})^{\top}A[\gamma]x^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\\ \left\langle A[\gamma]x+b[\gamma],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\gamma)\in{\cal G}(x,t)^{\perp}\end{array}\right\}.$
@@
Proposition 2.
Suppose Assumptions 1 and 2 hold. Let $(x,t)\in{\cal S}_{\textup{SDP}}$ where $2t=\sup_{\gamma\in\Gamma_{1}}[\gamma,q(x)]$. Let $(1,f)\in\operatorname{rint}({\cal F}(x,t))$. Then,
${\cal R}^{\prime}(x,t)=$
$\left\{(x^{\prime},t^{\prime})\in{\mathbb{R}}^{n+1}:\,\begin{array}[]{l}x^{\prime}\in\ker(A[f])\\ \left\langle A[\eta]x+b[\eta],x^{\prime}\right\rangle-t^{\prime}=0,\,\forall(1,\eta)\in{\cal G}(x,t)^{\perp}\end{array}\right\}.$
Proof.
Let $(1,f)\in\operatorname{rint}({\cal F}(x,t))$. By proposition 1, it suffices to prove $x^{\prime}\in\ker(A[f])$ if and only if
@@
3.4 Revisiting the setting of polyhedral $\Gamma$
Wang and Kılınç-Karzan [18] give sufficient conditions for convex hull exactness under the assumption that $\Gamma_{1}$ is polyhedral (equivalently, by remark 1, $\Gamma$ is polyhedral). This assumption holds, for example, when the set of quadratic forms $\left\{A_{\textup{obj}},A_{1},\dots,A_{m}\right\}$ is simultaneously diagonalizable. Specializing theorem 2 to this setting, we prove the following necessary and sufficient counterpart to [18, Theorem 1].
{restatable}